OpenStudy (anonymous):

Help ASAP! what is the GCF 0f 24a^3b and 36ab^2 thank u!!

OpenStudy (anonymous):

is it 12ab?

OpenStudy (jdoe0001):

\(\bf 24a^3b \quad and \quad 36ab^2\implies {\color{blue}{ 12}}\cdot 2{\color{blue}{ a}}a^2{\color{blue}{ b}} \quad and \quad {\color{blue}{ 12}}\cdot 3{\color{blue}{ ab}}\cdot b\)

OpenStudy (jdoe0001):

is it 12ab? \(\checkmark\)

OpenStudy (anonymous):

yay! thank u

OpenStudy (jdoe0001):

yw